OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of. y - 4 = 1/2(x - 1) y + 4 = 2(x - 1) y + 4 = 1/2(x - 1) y - 4 = 2(x + 1) Which of the following is the slope between the points (0, -10) and (-4, 0)? 2 - 5 2 - -- 5 5 - -- 2 5 - 2

myininaya (myininaya):

with the slope of what?

OpenStudy (anonymous):

im not sure. thats why i need help.

myininaya (myininaya):

Well then if you don't know the slope for the first one, no one can really help. The second one you do have enough info to find the slope.

OpenStudy (anonymous):

oh i thought u meant the second one, the answer choices for number 1 are the slopes.

myininaya (myininaya):

\[slope=m=\frac{ \Delta y}{\Delta x}=\frac{y_1-y_2}{x_1-x_2}\]

myininaya (myininaya):

\[(x_1,y_1)=(0,-10) \text{ and } (x_2,y_2)=(-4,0)\] Plug these numbers in Then use order of operations to simplify: \[\frac{y_1-y_2}{x_1-x_2}\]

myininaya (myininaya):

Like I said you need to know the slope of the first one. You haven't given this info. Read your first question again. I can help you with the second but you do not have enough info for the first question.

OpenStudy (anonymous):

oh im sorry, a slope of 1/2. i lookd it over and it didnt write. sorry

myininaya (myininaya):

Ok well we have started with second one so let's finish the second on and then we will go back to the first one. You need to enter in the values you have for x1,x2,y1,y2 into that formula above

OpenStudy (anonymous):

how do i do that ? sorry im not good at this at all.

myininaya (myininaya):

\[(x_1,y_1)=(0,-10) \text{ and } (x_2,y_2)=(-4,0) \]

myininaya (myininaya):

Do you see what I'm saying x1 is ? '''''''''''''''''' x2 is? '''''''''''''''''' y1 is? '''''''''''''''''' y2 is?

OpenStudy (anonymous):

okay i think

myininaya (myininaya):

You plug those numbers into this formula: \[\frac{y_1-y_2}{x_1-x_2}\]

myininaya (myininaya):

Replace y1 with -10 Replace y2 with -4 Replace x1 with 0 Replace x2 with -4 Then apply order of operations

OpenStudy (anonymous):

okay so it would be (0,-10)=(0,−10) and (-4,-4)=(−4,0)

OpenStudy (anonymous):

right ?

myininaya (myininaya):

No.... Ok \[\frac{y_1-y_2}{x_1-x_2}\] I told you to replace x1 with 0, x2 with -4 , y1 with -10, and y2 with -4 Can you replace them in the formula I just mentioned?

OpenStudy (anonymous):

im trying to. i see what you mean but i dont know how to write that formula in the box. but yes i can.

myininaya (myininaya):

Cool so what do you get when you do the subtraction on top? And then tell me what you got when you did the subtraction on bottom

OpenStudy (anonymous):

-6 on top and positive 4 on bottom.

myininaya (myininaya):

Yep! :) That is right! You did it. Good job. :)

myininaya (myininaya):

You can simplify -6/4 too.

OpenStudy (anonymous):

yay lol

myininaya (myininaya):

So I will let you reduce that on your own Let's go back to your first question

myininaya (myininaya):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2. y - 4 = 1/2(x - 1) y + 4 = 2(x - 1) y + 4 = 1/2(x - 1) y - 4 = 2(x + 1)

myininaya (myininaya):

Ok...So all these are in this form: \[y-y_1=m(x-x_1)\]

myininaya (myininaya):

This is called the point-slope form

myininaya (myininaya):

\[(x_1,y_1)=(1,-4), m=\frac{1}{2} \]

OpenStudy (anonymous):

ok

myininaya (myininaya):

You plugged those into that point-slope form and tell me what you get

myininaya (myininaya):

Replace that y1 with -4 that m with 1/2 that x1 with 1

OpenStudy (anonymous):

alright. i think i got it.

OpenStudy (anonymous):

1-4 --- 1−4 ========= 1/2=1/2

OpenStudy (anonymous):

right ?

myininaya (myininaya):

You are pluggin' into this formula: \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

i did

myininaya (myininaya):

You replaced y1 with -4 m with 1/2 and x1 with 1 That is all you need to do You don't need any extra work after that really